Can you help me solve this?
radical sign x-1 = x-3
The radical sign is only by x-1
Thanks for the help!
CH

Answer by Edwin McCravy(20056)

Can you help me solve this?
radical sign x-1 = x-3
The radical sign is only by x-1

 ___
√x-1 = x-3

Square both sides

  _____
(√x - 1)² = (x - 3)²

When you square a square root, you 
take away the radical

    x - 1 = (x - 3)²

Now you square the right side by writing
(x - 3)² as (x - 3)(x - 3)

    x - 1 = (x - 3)(x - 3)

You finish it! It will become a quadratic 

equation.  You will get solutions 2 and 5

However you must ALWAYS check radical equations! 

You will find that the 2 does not check in the

original equation, but the 5 does check.

So you discard the 2 solution as an extraneous 

solution, and the only correct solution is 5.

Edwin